### v1

# Bimanual Regrasp Planning and Control for Eliminating Object Pose Uncertainty

Abstract
Precisely grasping an object is a challenging task due to pose uncertainties. Conventional methods have used cameras and fixtures to reduce object uncertainty. They are effective but require intensive preparation, such as designing jigs based on the object geometry and calibrating cameras with high-precision tools fabricated using lasers. In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera. Our method is based on the concepts that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact. Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to eliminate uncertainties in the object pose. We evaluated the proposed method on different initial object uncertainties and verified that the method have satisfactory repeatability accuracy. It outperforms an AR marker detection method implemented using cameras and laser jet printers under standard laboratory conditions.
Index Terms:
Grasping Uncertainty, Bimanual Manipulation
I Introduction
Asignificant challenge in robotic manipulation lies in addressing the uncertainties associated with object grasping. The uncertainties often arise from errors in environmental registration, inaccuracies in object pose recognition, and unbalanced contact during grasping that leads to pose deviations. The uncertainties can result in discrepancies between the actual and expected pose of objects or tools, potentially causing task failures. Existing methods used to address uncertainty problems typically fall into two categories: Sensor-based correction methods [1][2] and fixture-based constraint methods [3][4]. Sensor-based methods leverage additional hardware, such as force sensors or tactile sensors, to detect and mitigate uncertainties. Fixture-based methods, on the other hand, rely on structuring the environment or using special-purpose structures to constrain object pose. However, both approaches have notable limitations, as they often require additional hardware adjustments or modifications to the environment.
Figure 1: (a) A workpiece is difficult to be accurately grasped without fixture-based positioning. (b) Flat finger pads of parallel grippers can help reduce uncertainty in the hand opening/closing direction. After grasping, the object’s uncertainty is constrained within the gray plane shown in the subfigure. (c) Three orthogonal grasps by parallel grippers constrain an exact pose. This paper proposes a bimanual regrasp planning method that generates regrasping motion sequence by consecutively executing three orthogonal grasps to eliminate object pose and grasp uncertainty.
This paper proposes a novel bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to eliminate uncertainties. The proposed approach does not require external fixtures or additional sensors, making it more flexible for diverse robotic manipulation scenarios. The fundamental concepts of the proposed method are illustrated in Fig. 1(b) and (c). It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact. Three such grasps collectively constrain the object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triples to simultaneously eliminate uncertainties in the object pose. However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage. To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose. By integrating regrasp planning with admittance control, our method iteratively eliminates uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
In the experiments, we implemented the proposed method and evaluated its performance on different initial object uncertainties. The results demonstrate that the method achieves satisfactory accuracy without the need for external fixtures or additional sensors. Furthermore, we compared the accuracy of the proposed method with that of AR-based detection. The results indicate that our method outperforms the AR marker detection implemented using cameras and laser jet printers under standard laboratory conditions. The proposed approach leverages the robot’s inherent geometric constraints and force sensors to eliminate uncertainties in the grasped object. It improves task success rates by addressing uncertainties in a systematic and hardware-efficient manner.
II Related Work
The primarily related work of this paper focuses on regrasp planning and the elimination of grasping errors. Regrasp planning aims at automatically generating grasp sequences to accomplish object manipulation. It involves two main categories: One aims to identify discrete pick-and-place sequences. For example, Liu et al. [7] developed regrasp planning using multiple mobile robots. The method achieves flipping manipulation with the help of regrasping. Kim et al. [8] developed a randomized planning method for regrasp planning. Baek et al. [9] considered non-prehensile grasping in regrasp planning and secured power grasp for stable holding. Qin et al. [10] presented a regrasp planning method for installing a deformable linear object. Rigidity constraints were considered when planning regrasp motion. Mitrano et al. [11] studied a similar problem and proposed considering avoiding environment-grasp-body loops to accelerate grasp selection and planning. Sundaralingam et al. [12] extended regrasp planning to multi-fingered hand. Murooka et al. [13] proposed the foot-step and regrasping planning method for a humanoid robot to simultaneously determine foot steps and hand switch. The other focuses on continuous adjustments. For instance, Maximo et al. [14] presented the grasp space method for continuous regrasp (rolling or sliding) that maintains force closures. Hu et al. [15] formulated regrasp planning as a multi-modal problem, where three manifolds, including the regrasp manifold, the transfer manifold, and the slide manifold, were planned across to determine action sequences and thus plan motion. Cheng et al. [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds. The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
One of the most critical challenges in regrasp planning is the planning speed. Researchers have explored various methods to enhance efficiency. For example, Cho et al. [17] developed a fast regrasp planning algorithm using look-up tables and two-finger parallel grippers. Chiu et al. [18] proposed a reinforcement learning method to accelerate the planning of regrasp poses. Their actions were defined in the local frame of a robot hand, thereby resolved the dependencies on robotic kinematics. Wada et al. [19] significantly improved planning speed by using a neural network to filter waypoint poses and quickly select grasp poses. Xu et al. [20] proposed learning reorientation object poses and then plan regrasp motion while considering the learned middle reorientation poses. Learning-based methods have demonstrated effectiveness in improving regrasp planning efficiency. While there are other numerous publications, this paper does not focus heavily on this aspect and thus will not delve into further details.
Another critical challenge in regrasp planning is uncertainty. Repeated regrasping can result in error accumulation, causing the grasp pose to deviate significantly from the planned outcome. The deviation can prevent the robot from successfully executing subsequent tasks. To address the uncertainty problems, Bauza et al. [21] integrated tactile information into regrasp planning and achieved precise and robust pick and placement for regrasping. Bronars et al. [22] proposed a simultaneous tactile estimation and control method to continuously plan sliding regrasp motion while maintaining grasping precisions. Gualtieri et al. [23] estimated the quality of grasp candidates against uncertainty based on point cloud completion, and optimized regrasp sequences to be robust. Zhang et al. [24] used reinforcement learning to learn a regrasp policy that optimizes grasp stability. Both vision and visual, tactile information were used for learning. This study can be partially classified as research aimed at reducing uncertainties in grasping poses during the regrasp planning process. However, our primary goal is not to achieve precise posse estimation for regrasping. Instead, we focus on actively leveraging the consecutive geometric constraints inherent in the regrasping process [25][26][27][28] to eliminate errors.
III Regrasp Planning Using Orthogonal Triples
The essence of using regrasp to eliminate uncertainty lies in leveraging the parallel finger pads of a gripper to constrain an object’s pose to a plane. By performing three successive regrasp using parallel-finger grippers along mutually orthogonal directions, an object can be constrained to three mutually orthogonal planes, and thus be precisely located. Therefore, the first step to achieve the goal of this study is to identify grasp poses along three mutually orthogonal directions. We are going to introduce the grasp planning and grouping methods in the first subsection below. Once the grasp poses and groups are determined, search and motion planning techniques can be employed to generate regrasp actions and consequently the robot motion required to eliminate uncertainties. The second subsection below details the specific methods for search and motion planning.
III-A Orthogonal Grasp Triples
[…]
Figure 2: (a) Antipodal grasp planning. (b) Candidate grasp poses planned using the antipodal method. (c) Separating the candidate grasp poses into three mutually orthogonal groups considering their opening and closing directions. Each color in the subfigure represents one group.
Subsequently, based on the results of the grasp planning, we proceed by separating them into groups. Since our goal is to identify three mutually orthogonal grasp poses, we set our grouping criterion to be based on the opening and closing directions of the gripper fingers. Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 1 Note 1 Note 1In this context, the direction of $\boldsymbol{v}_{i}$ is not considered. That is, $\boldsymbol{v}_{i}$ assumed identical to $-\boldsymbol{v}_{i}$.. Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …The candidate grasp poses in Fig. 2(b) can be right classified into three mutually orthogonal groups. They are visualized in Fig. 2(c) using red, green, and blue colors.
After obtaining the grasp groups, we perform triple combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint. To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triple combination.
$\mathtt{score}=|\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{j}|+|\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{k}|+|\boldsymbol{v}_{j}\cdot\boldsymbol{v}_{k}|$ (1)
The $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ in the above equation represent the opening and closing directions of three different groups in a triple. We sort the scores of all triples and prioritize them with a score closest to 0 for regrasp planning and uncertainty elimination. The closer the score value is to 00, the more orthogonal the three grasps from the groups are to each other. By performing successive regrasp using the prioritized triples, object uncertainties can be effectively minimized. The example shown in Fig. 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero. In more general cases, there may be multiple candidate groups and triple combinations. The triples may have non-zero scores and regrasp should be performed by prioritizing the ones with smaller scores.
III-B Regrasp Planning
For the sorted triples, motion planning is sequentially performed on each of them, and the first successful solution is output as a valid candidate. Note that each triple comprises three groups of grasp poses. During planning, one pose from each group needs to be selected to form a combination of three grasps for computation. In this approach, we do not explicitly extract the three-grasp combinations as a separate step. Instead, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. The grasps in each group are sequentially assigned to the two robotic arms for planning. In this way, we incrementally explore the combinations of the grasp triples, and output successful motion once found. The detailed algorithmic flow is illustrated in Fig. 3. Note that the process illustrated in the figure is merely an incremental search. If a large amount of data can be accumulated using the proposed method in the future, we strongly recommend exploring learning-based methods to improve efficiency.
Figure 3: Regrasp planning used in this study is a simple incremental search method, which sequentially traverses grouped grasp triples until a valid solution is found. Since the regrasp planning algorithm itself is not the primary focus of this paper, further optimization aspects such as acceleration or grasp selection refinement are not explored.
IV Conforming to Object Pose Uncertainty
After planning a feasible sequence of bimanual robot motions, the next step is execution. However, due to uncertainties in the object’s pose, a significant discrepancy may exist between the real-world pose and the planning simulation, making pure position control infeasible. This subsection discusses this limitation and highlights the necessity of conforming to object pose uncertainty. It then introduces the admittance control method and explains how it enables the system to adapt to uncertain object poses during handover regrasping.
Consider the scenario shown in Fig. 4(a), where the left hand has already moved the object from its initial pose to the handover pose, and the right hand is approaching to perform the regrasp. The target pose of the right hand is planned based on an assumed pose in the simulation. However, due to uncertainties in the real world, the object’s assumed pose in the simulation deviates from its actual pose, as illustrated by the orange dashed frames and red solid frames in Fig. 4(a). Consequently, the right hand’s grasping pose becomes misaligned with the actual object pose. Applying position control for grasping in this scenario may lead to undesirable outcomes. In mild cases, as shown in Fig. 4(b) and (c), the object may undergo deformation. In more severe cases, it may sustain damage. Therefore, rigidly executing the planned grasping motion solely with position control is not advisable.
Figure 4: (a) The right hand approaches to grasp the object. Due to uncertainties, the object is not positioned exactly at the center of the two fingers as assumed but is instead slightly shifted upward. (b) At this point, if position control is used, the upper-side finger will make contact with the object first, exerting excessive force on it and causing deformation. (c) Comparison of object deformation before and after grasping.
To address this issue, this study employs admittance control during robotic grasping and enables the manipulator’s pose to conform to the object’s pose. The basic concept of admittance control is to establish a mass-damper-spring system’s motion equation based on the external force exerted on the robot and the desired target force . The robot then follows the solution of this equation to adjust its motion accordingly, thereby accommodating external forces. For a desired robot position $x_{d}$, the admittance control equation is
$M(\ddot{\mathit{x}}_{p}-\ddot{\mathit{x}}_{d})+B(\dot{\mathit{x}}_{p}-\dot{\mathit{x}}_{d})+K(\mathit{x}_{p}-\mathit{x}_{d})=\mathit{f}+\mathit{f}_{d}.$ (2)
Here, $x_{p}$, $\dot{x_{p}}$, and $\ddot{x_{p}}$ are respectively the measured current robot end-effector position, velocity, and acceleration. $M$, $B$, $K$ denote the virtual inertia, damping, and stiffness parameters. $f$ is the force applied to the system, while $f_{d}$ represents the target force exerted by the robot after applying admittance control.
In our specific grasping scenario, the target force is set to zero, as our objective is to ensure that the robot does not exert any external force on the object beyond the necessary grasping force, thereby preventing unintended movement, deformation, or damage. During the grasping process, when one of the robotic fingers first makes contact with the object, it receives a reaction force from the object, which corresponds to the external force $f$ in equation (2). Based on the robot’s current $x_{p}$, $\dot{x_{p}}$, $\ddot{x_{p}}$, as well as the force $f$ measured using a force/torque sensor installed at the robot manipulator’s wrist, we can compute the desired position $x_{d}$, velocity $\dot{x_{d}}$, and acceleration $\ddot{x_{d}}$ using the equation. These computed values are then used to control the robot, enabling it to conform to the object’s uncertain pose during the grasping process.
Fig. 5 illustrates the effect of applying admittance control to the grasping process shown in Fig. 4. In Fig. 5(a), the upper-side finger first makes contact with the object and experiences a reaction force $f$. The admittance control then adjusts the robotic hand by moving it in a direction that balances this force, as shown in Fig. 5(b). The adjustment continues until both fingers are in contact with the object and their grasping forces reach equilibrium. At this point, $f$ disappears and the grasping forces become internal forces within the whole gripper-object system. The final grasping result is depicted in Fig. 4(c). Unlike position control, admittance control ensures that the object’s pose remains unchanged after grasping.
Figure 5: Admittance control helps avoid deformation or damage. (a) When the upper-side finger makes contact with the object, a reaction force $f$ is detected. (b) Admittance control adjusts the robot hand by moving it in a direction that balances $f$. (c) After admittance control, the object is firmly grasped, and no deformation occurs in its shape.
V Pose Estimation Using Conformed Grasp Poses
Next, we present the method for estimating the object’s pose. In the planning simulation, we used an assumed object pose. This pose serves as the basis for all grasping and planning processes. Since there is no uncertainty in the simulation, the final object pose remains deterministic. Here, we denote the object’s final pose in the planning simulation as $\boldsymbol{p}_{o_{s}}$ and $\boldsymbol{R}_{o_{s}}$, and denote the three consecutive grasping poses of the robot in the simulation as $\boldsymbol{g}_{s_{1}}$, $\boldsymbol{g}_{s_{2}}$, and $\boldsymbol{g}_{s_{3}}$. Each $\boldsymbol{g}_{s_{i}}$ consists of ($\boldsymbol{p}_{s_{i}}$, $\boldsymbol{R}_{s_{i}}$, $d_{s_{i}}$), where $\boldsymbol{p}_{s_{i}}$ and $\boldsymbol{R}_{s_{i}}$ represent the position and rotation of the grasping center point, and $d_{s_{i}}$ denotes the jaw opening width. On the other hand, by executing the planned motion in the real world, we obtain three consecutive grasping poses of the bimanual robot in the real world as $\boldsymbol{g}_{r_{1}}$, $\boldsymbol{g}_{r_{2}}$, and $\boldsymbol{g}_{r_{3}}$. Each $\boldsymbol{g}_{r_{i}}$ similarly consists of ($\boldsymbol{p}_{r_{i}}$, $\boldsymbol{R}_{r_{i}}$, $d_{r_{i}}$). The first grasping pose $\boldsymbol{g}_{r_{1}}$ aligns exactly with the planned result, while the second grasping pose $\boldsymbol{g}_{r_{2}}$ and the third grasping pose, $\boldsymbol{g}_{r_{3}}$, are adjusted through admittance control to conform to the object’s actual pose. They are different from $\boldsymbol{g}_{s_{2}}$ and $\boldsymbol{g}_{s_{3}}$. Our objective is to estimate the real-world object pose $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ based on these known values.
To address this issue, let us first consider the simulation scenario. Since our three grasping poses involve planar contacts and are mutually orthogonal, the object constrained by them should have a unique pose. The pose can be computed as follows. For each grasp, we can obtain a corresponding opening and closing vector $\boldsymbol{v}_{i}=\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$. Here, $\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$ denotes the $\boldsymbol{y}$ axis (2nd column) of the rotation matrix. We assume the robot hand’s opening and closing direction is aligned with its local $\boldsymbol{y}$ axis. Since the three grasping poses are mutually orthogonal, they uniquely determine the following rotation matrix
$\boldsymbol{R}_{g_{s}}=\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})&\boldsymbol{R}_{s_{2}}(\boldsymbol{y})&\boldsymbol{R}_{s_{3}}(\boldsymbol{y})\end{bmatrix}.$ (3)
Meanwhile, for each grasp, its two parallel fingers define a midplane. Given that the three grasps are mutually orthogonal, their corresponding midplanes intersect at a unique intersection point. This intersection represents a unique position constraint imposed by the grasping configuration. It can be computed using
$\boldsymbol{p}_{g_{s}}=(\boldsymbol{R}_{g_{s}}^{\top})^{-1}\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{1}}}\\ \boldsymbol{R}_{s_{2}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{2}}}\\ \boldsymbol{R}_{s_{3}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{3}}}\end{bmatrix}$ (4)
The ($\boldsymbol{p}_{g_{s}}$, $\boldsymbol{R}_{g_{s}}$) obtained from the above two equations are the unique pose determined by the three grasping poses in the planning simulation. Note that it is not the object’s final pose ($\boldsymbol{p}_{o_{s}}$, $\boldsymbol{R}_{o_{s}}$). However, since both are uniquely determined. There is a unique offset between them written as
$\boldsymbol{p}_{\text{diff}}=\boldsymbol{p}_{o_{s}}-\boldsymbol{p}_{g_{s}},~{}\boldsymbol{R}_{\text{diff}}=\boldsymbol{R}_{g_{s}}^{\top}\boldsymbol{R}_{o_{s}}$ (5)
Fig. 6 illustrates the three planes formed by three grasping poses and their intersection point.
Figure 6: Three orthogonal grasping poses with planar contacts determine three planes that intersect at a unique point. Thus, a unique position $\boldsymbol{p}_{g_{s}}$ is determined. Meanwhile, the opening-closing directions of the three gasping poses help constrain a unique rotation $\boldsymbol{R}_{g_{s}}$
In the real world, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ are uniquely constrained by the grasping poses and have the same offset from the pose determined by the grasping poses. That is
$\boldsymbol{p}_{o_{r}}=\boldsymbol{p}_{g_{r}}+\boldsymbol{p}_{\text{diff}},~{}\boldsymbol{R}_{o_{r}}=\boldsymbol{R}_{g_{r}}\boldsymbol{R}_{\text{diff}}.$ (6)
Thus, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ can be computed. Here, $\boldsymbol{p}_{g_{r}}$ and $\boldsymbol{R}_{g_{r}}$ can be obtained similarly to equations (4) and (3) using ($\boldsymbol{p}_{r_{1}}$, $\boldsymbol{R}_{r_{1}}$), ($\boldsymbol{p}_{r_{2}}$, $\boldsymbol{R}_{r_{2}}$), and ($\boldsymbol{p}_{r_{3}}$, $\boldsymbol{R}_{r_{3}}$).
VI Experiments and Analysis
This section presents the real-world validation of the proposed method. Our experimental platform consists of two UR3e robots from Universal Robots. Their installation setup is shown in Fig. 7. The UR3e series robots are equipped with high-precision force/torque sensors at their end effectors, which facilitate the implementation of admittance control. Each robot is installed with a Robotiq Hand-E gripper from Robotiq. The opening range of the gripper is 0$\sim$50 mm. The regrasp planning is performed using the WRS simulation environment Note 2 Note 2 Note 2https://github.com/wanweiwei07/wrs. Fig. 6 in the previous section provided a view of the simulation environment. The coordinate system of our experimental platform is defined with its origin at the intersection of the central plane between the two robot arms and the rear edge of the table, as illustrated by the colored frame in Fig. 7. The $\boldsymbol{x}$-axis points forward, the $\boldsymbol{y}$-axis points towards the left arm, and the $\boldsymbol{z}$-axis aligns with the normal direction of the tabletop surface.
[…]
To ensure clarity in comparison and analysis, we used the simple L-shaped object shown in Fig. 7 and assumed a single initial pose in the planning simulation. The object has a long edge of $125$ mm and a short edge of $100$ mm, with a square cross-section of $25$ mm $\times$ $25$ mm. The surfaces of the L-shaped object are mutually perpendicular. They could align with the three axes of the coordinate system and facilitate analysis. The initial pose of the object in the simulation was set to $(700,300,1.25)$ mm. Its rotation was set to be aligned with the coordinate frame of the robot platform (recall the colored frame shown in Fig. 7). The long edge aligns with the $\boldsymbol{y}$ axis. The short edge aligns with the $\boldsymbol{z}$ axis. Similarly to the initial pose, we set the regrasp pose of the object in the simulation to be at a unique fixed location $(680,50.0,120.0)$ mm, with the rotation identical to the initial one to facilitate analysis. The location is in the middle of the two arms, making it easy to plan effective bimanual regrasp motions. The regrasp motion planned based on the parameter settings and the proposed planning methods led to orthogonal triples where the first selected grasp is along the $\boldsymbol{y}$ axis, the second grasp is along the $\boldsymbol{z}$ axis, and the third grasp is along the $x$ axis. Together, the three grasps uniquely constrain the object’s pose.
In the real world, we manually introduced uncertainties in multiple directions to the object for experimental evaluation. These uncertainties are illustrated in Fig. 8. Fig. 8(a) represents initial object poses obtained by introducing uncertainties in the horizontal directions (back, front, left, and right) relative to a center pose shown in Fig. 8(a.v). The center pose closely corresponds to the initial pose of the object used in the simulation. Note that the center pose is not exactly the initial pose in the simulation. The front, back, left, and right movements are approximate ±$\pm 12.5$ mm and are also not exact. We are not enforcing precise placement of the object when preparing these poses. The reason is that preparing exact uncertainty in the real world can be challenging. Finger contact during the grasping process naturally causes additional errors. Specifying exact uncertainty values and evaluating absolute accuracy would not provide meaningful insights. We analyze repeatability errors rather than absolute errors to evaluate the effectiveness of our method in eliminating uncertainties. In addition to the horizontal uncertainties, we also prepared initial poses with uncertainties in the upward, upper-front, and upper-back directions. These poses are shown in Fig. 8(b). The L-shape object is elevated using a white block while being shifted back and forward with the same approximate ±$\pm 12.5$ mm offset when preparing these uncertain poses.
Figure 8: (a) $9$ uncertain poses in the horizontal directions. (a.i) BL – Back Left. (a.ii) BC – Back Center. (a.iii) BR – Back Right. (a.iv) L – Left. (a.v) C – Center. (a.vi) R – Right. (a.vii) FL – Front Left. (a.viii) FC – Front Center. (a.ix) FR – Front Right. (b) $3$ uncertain poses in upward directions. (b.i) DC – Down Center (for reference). (b.ii) UC – Upper Center. (b.iii) Upper Back. (b.iv) Upper Front.
Using the planned robot motions, we performed regrasping and applied admittance control following the proposed method for each initial object pose in Figs. 8(a) and (b), and estimated their final object poses using equation 6. The estimation results are summarized in the “Proposed method” columns of Table I. The results corresponding to (a) and (b) are separated by a double solid line. From the results, we can observe that the estimation performs reasonably well. In the case of the uncertain object placement shown in Figs. 8(a), the estimated $\boldsymbol{p}_{o_{r}}$ vector has a mean and variance of $(9.82,0.002)$ mm and $(898.33,0.019)$ mm in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions respectively. The low variance aligns well with our intuition – For the uncertain poses listed in Fig. 8(a), the object is placed on the table. Its position in the $\boldsymbol{z}$-direction remains the same. Meanwhile, the object is always grasped by fingers closing from the $\boldsymbol{y}$-direction. Its uncertainty in the $\boldsymbol{y}$-direction is constrained first grasp’s two fingers. Therefore, the final estimated poses should be consistent (have a low variance) in these two directions. On the other hand, the first values of the estimated $\boldsymbol{p}_{o_{r}}$ can be clearly categorized into three distinct groups. Among them, BC and BR form one group, with a mean and variance of $(687.05,2.1)$ mm. L, C, and R constitute another group, with a mean and variance of $(699.3,1.4)$ mm. Finally, FL and FC belong to the third group, with a mean and variance of $(709.45,2.1)$ mm. This result is also consistent with our initial placement. The object displacement in the $\boldsymbol{x}$ direction can be clearly distinguished, and the repeatability is satisfying. The variance of each distinct group falls within the range of the manual placement uncertainty. Similar analysis and conclusions can be drawn from the results for the uncertain object placements shown in Figs. 8(b). They are in the UB, UC, and UF rows. The estimated $\boldsymbol{p}_{o_{r}}$ vectors have nearly the same values in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions and clearly distinct values in the $\boldsymbol{x}$ direction. For rotation, all estimated results are generally aligned with the coordinate axes. They have an average deviation of $0.59^{\circ}$ with $0.11^{\circ}$ variance from rotation matrix $\begin{bmatrix}\boldsymbol{x}&\boldsymbol{z}&-\boldsymbol{y}\end{bmatrix}$. This outcome is consistent with the fact that the object’s initial placement is always on a plane and that the opening-closing direction of the first grasp is along the $\boldsymbol{y}$ axis. Its repeatability is also satisfying considering the small variance. Note that the BL and FR rows in the table do not contain values. This is because the object’s pose uncertainty was too large, causing the robot to collide with the object during execution. The maximum tolerable uncertainty in our method is constrained by the gripper’s opening width. When the uncertainty exceeds half of the gripper’s opening width, collisions may occur and lead to execution failure. Additionally, the table does not include estimation results for DC pose. This pose is only provided to give readers a reference for comparison before the object is elevated, and thus, no estimation was performed for it.
TABLE I: Pose estimation results
Type Exec. Proposed method AR marker
$\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$ $\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$
BL Fail. - - - -
BC Succ. $(685.6,49.8,118.2)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.01&0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(715.4,56.5,128.3)$ $\begin{bmatrix}1.00&-0.01&0.07\\ 0.07&-0.02&-1.00\\ 0.01&1.00&-0.02\end{bmatrix}$
BR Succ. $(688.5,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.02&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(719.5,34.8,146.3)$ $\begin{bmatrix}1.00&-0.06&0.00\\ 0.00&-0.02&-1.00\\ 0.06&1.00&-0.02\end{bmatrix}$
L Succ. $(699.5,49.9,118.5)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(740.2,54.1,131.5)$ $\begin{bmatrix}1.00&-0.03&0.05\\ 0.05&-0.01&-1.00\\ 0.03&1.00&-0.00\end{bmatrix}$
C Succ. $(697.8,49.9,118.1)$ $\begin{bmatrix}1.00&-0.00&0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(726.9,36.8,144.5)$ $\begin{bmatrix}1.00&-0.07&-0.01\\ -0.01&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
R Succ. $(700.7,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(726.3,55.6,145.8)$ $\begin{bmatrix}1.00&-0.07&0.07\\ 0.06&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
FL Succ. $(710.9,49.8,118.4)$ $\begin{bmatrix}1.00&0.01&-0.00\\ 0.03&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(741.3,57.7,128.4)$ $\begin{bmatrix}0.99&-0.01&0.10\\ 0.10&-0.01&-0.99\\ 0.01&1.00&-0.01\end{bmatrix}$
FC Succ. $(708.0,49.8,118.5)$ $\begin{bmatrix}1.00&0.00&-0.00\\ 0.02&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(756.1,29.4,110.6)$ $\begin{bmatrix}1.00&-0.00&0.01\\ 0.01&-0.03&-1.00\\ 0.00&1.00&-0.03\end{bmatrix}$
FR Fail. - - - -
UB Succ. $(689.8,49.8,131.8)$ $\begin{bmatrix}1.00&0.01&0.00\\ -0.02&-0.01&-1.00\\ -0.03&1.00&0.00\end{bmatrix}$ $(717.5,38.3,157.9)$ $\begin{bmatrix}1.00&-0.06&-0.02\\ -0.02&-0.03&-1.00\\ 0.06&1.00&-0.03\end{bmatrix}$
UC Succ. $(698.7,49.8,131.6)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(732.4,32.1,160.6)$ $\begin{bmatrix}1.00&-0.07&0.02\\ 0.02&-0.02&-1.00\\ 0.07&1.00&-0.02\end{bmatrix}$
UF Succ. $(706.6,49.8,131.6)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ 0.01&-0.02&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(735.6,36.6,158.5)$ $\begin{bmatrix}1.00&-0.08&0.02\\ 0.01&-0.02&-1.00\\ 0.08&1.00&-0.02\end{bmatrix}$
[…]
VII Conclusions and Future Work
This paper proposes a method for eliminating object uncertainties by leveraging the grasping constraints of parallel grippers with flat finger pads and a bimanual regrasp planning method. We applied this method to estimate grasping errors for a roughly placed L-shaped object. The results demonstrate that the proposed approach effectively eliminates uncertainties and achieves high repeatability accuracy.
Currently, this method is only applicable when all three grasps involve planar contact. For curved surface contacts, additional grasps may be required to fully constrain uncertainties. Further exploring the necessary grasp combinations will be an important future topic.

### v3

@@
-Bimanual Regrasp Planning and Control for Eliminating Object Pose Uncertainty
+Bimanual Regrasp Planning and Control for Active Reduction of Object Pose Uncertainty
 Abstract
@@
 In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera.
-Our method is based on the concepts that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact.
+Our method is based on the concept that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact.
 Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state.
-Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to eliminate uncertainties in the object pose.
-We evaluated the proposed method on different initial object uncertainties and verified that the method have satisfactory repeatability accuracy.
-It outperforms an AR marker detection method implemented using cameras and laser jet printers under standard laboratory conditions.
+Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to actively reduce uncertainties in the object pose.
+We evaluated the proposed method on different initial object uncertainties and verified that it had good repeatability.
+The deviation levels of the experimental trials were on the same order of magnitude as those of an optical tracking system, demonstrating strong relative inference performance.
 Index Terms:
-Grasping Uncertainty, Bimanual Manipulation
+Grasping uncertainty, Regrasping, Bimanual manipulation
 I Introduction
@@
 The uncertainties often arise from errors in environmental registration, inaccuracies in object pose recognition, and unbalanced contact during grasping that leads to pose deviations.
-The uncertainties can result in discrepancies between the actual and expected pose of objects or tools, potentially causing task failures.
+The uncertainties can result in discrepancies between the actual and expected poses of objects or tools, potentially causing task failures.
 Existing methods used to address uncertainty problems typically fall into two categories: Sensor-based correction methods [1][2] and fixture-based constraint methods [3][4].
@@
 Fixture-based methods, on the other hand, rely on structuring the environment or using special-purpose structures to constrain object pose.
-However, both approaches have notable limitations, as they often require additional hardware adjustments or modifications to the environment.
-Figure 1: (a) A workpiece is difficult to be accurately grasped without fixture-based positioning.
-(b) Flat finger pads of parallel grippers can help reduce uncertainty in the hand opening/closing direction.
-After grasping, the object’s uncertainty is constrained within the gray plane shown in the subfigure.
-(c) Three orthogonal grasps by parallel grippers constrain an exact pose.
-This paper proposes a bimanual regrasp planning method that generates regrasping motion sequence by consecutively executing three orthogonal grasps to eliminate object pose and grasp uncertainty.
-This paper proposes a novel bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to eliminate uncertainties.
-The proposed approach does not require external fixtures or additional sensors, making it more flexible for diverse robotic manipulation scenarios.
-The fundamental concepts of the proposed method are illustrated in Fig.
+However, both approaches have notable limitations, as they often require additional hardware adjustments or environment modifications.
+Figure 1: (a) Precisely grasping an object is difficult.
+(a.i) Before grasping.
+(a.ii) After grasping.
+(a.iii) Comparison.
+(b) Flat finger pads of a parallel gripper help constrain uncertainty to a plane.
+(c) Three orthogonal parallel grasps constrain three planes, leading to an exact pose.
+This paper proposes a bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to reduce uncertainties.
+It does not require external fixtures or additional sensors and is more flexible for diverse robotic manipulation scenarios.
+The fundamental concept of the proposed method is illustrated in Fig.
 1(b) and (c).
 It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact.
-Three such grasps collectively constrain the object’s position and orientation to a unique state.
-Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triples to simultaneously eliminate uncertainties in the object pose.
+Three such grasps collectively constrain the object’s position and orientation to a unique state Note 1 Note 1 Note 1This unique pose constraint holds under the following assumptions: i) The object is sufficiently rigid such that the grasping force is transmitted and sensed through the force sensors, rather than causing deformation.
+ii) The contact between the gripper fingers and the object is assumed to be flat surface contact with sufficient friction to prevent slippage.
+iii) The maximum uncertainty in the object pose is within the gripper’s opening range, ensuring valid contact during grasping..
+Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triplets to simultaneously reduce uncertainties in the object pose.
 However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage.
 To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose.
-By integrating regrasp planning with admittance control, our method iteratively eliminates uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
+By integrating regrasp planning with admittance control, our method iteratively reduces uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
 In the experiments, we implemented the proposed method and evaluated its performance on different initial object uncertainties.
-The results demonstrate that the method achieves satisfactory accuracy without the need for external fixtures or additional sensors.
-Furthermore, we compared the accuracy of the proposed method with that of AR-based detection.
-The results indicate that our method outperforms the AR marker detection implemented using cameras and laser jet printers under standard laboratory conditions.
-The proposed approach leverages the robot’s inherent geometric constraints and force sensors to eliminate uncertainties in the grasped object.
-It improves task success rates by addressing uncertainties in a systematic and hardware-efficient manner.
+The results demonstrate that the method achieves satisfactory repeatability.
+The deviation levels of remaining uncertainties were the same order of magnitude as those of an optical tracking system, demonstrating that the method provides strong relative inference performance.
+By leveraging the robot’s inherent geometric constraints and force sensors, the proposed approach can help actively reduce uncertainties in the grasped object.
 II Related Work
-The primarily related work of this paper focuses on regrasp planning and the elimination of grasping errors.
-Regrasp planning aims at automatically generating grasp sequences to accomplish object manipulation.
-It involves two main categories: One aims to identify discrete pick-and-place sequences.
+II-A Categories of Regrasp Planning Methods
+Regrasp planning aims to automatically generate grasp sequences to accomplish object manipulation. It involves two main categories: discrete and continuous.
+Discrete regrasp planning focuses on identifying discrete pick-and-place sequences.
 For example, Liu et al.
@@
 Baek et al.
-[9] considered non-prehensile grasping in regrasp planning and secured power grasp for stable holding.
+[9] considered non-prehensile grasping in regrasp planning and secured power grasps for stable holding.
 Qin et al.
@@
 Sundaralingam et al.
-[12] extended regrasp planning to multi-fingered hand.
+[12] extended regrasp planning to a multi-fingered hand.
 Murooka et al.
-[13] proposed the foot-step and regrasping planning method for a humanoid robot to simultaneously determine foot steps and hand switch.
-The other focuses on continuous adjustments.
+[13] proposed the foot-step and regrasping planning method for a humanoid robot to simultaneously determine foot steps and hand switches.
+Continuous regrasp methods, on the other hand, aim to adjust object poses through controlled rolling or sliding motions while maintaining contact.
 For instance, Maximo et al.
@@
 The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
+II-B Key Challenges and the Focus of this Work
 One of the most critical challenges in regrasp planning is the planning speed.
@@
 [18] proposed a reinforcement learning method to accelerate the planning of regrasp poses.
-Their actions were defined in the local frame of a robot hand, thereby resolved the dependencies on robotic kinematics.
+Their actions were defined in the local frame of a robot hand, thereby resolving the dependencies on robotic kinematics.
 Wada et al.
@@
 Xu et al.
-[20] proposed learning reorientation object poses and then plan regrasp motion while considering the learned middle reorientation poses.
+[20] proposed learning reorientation object poses and then planning regrasp motion while considering the learned middle reorientation poses.
 Learning-based methods have demonstrated effectiveness in improving regrasp planning efficiency.
-While there are other numerous publications, this paper does not focus heavily on this aspect and thus will not delve into further details.
+While there are numerous other publications, this paper does not focus heavily on this aspect and thus will not delve into further details.
 Another critical challenge in regrasp planning is uncertainty.
@@
 Bronars et al.
-[22] proposed a simultaneous tactile estimation and control method to continuously plan sliding regrasp motion while maintaining grasping precisions.
+[22] proposed a simultaneous tactile estimation and control method to plan sliding regrasp motion while maintaining grasping precision continuously.
 Gualtieri et al.
-[23] estimated the quality of grasp candidates against uncertainty based on point cloud completion, and optimized regrasp sequences to be robust.
+[23] estimated the quality of grasp candidates against uncertainty based on point cloud completion and optimized regrasp sequences to be robust.
 Zhang et al.
 [24] used reinforcement learning to learn a regrasp policy that optimizes grasp stability.
-Both vision and visual, tactile information were used for learning.
-This study can be partially classified as research aimed at reducing uncertainties in grasping poses during the regrasp planning process.
-However, our primary goal is not to achieve precise posse estimation for regrasping.
-Instead, we focus on actively leveraging the consecutive geometric constraints inherent in the regrasping process [25][26][27][28] to eliminate errors.
-III Regrasp Planning Using Orthogonal Triples
-The essence of using regrasp to eliminate uncertainty lies in leveraging the parallel finger pads of a gripper to constrain an object’s pose to a plane.
-By performing three successive regrasp using parallel-finger grippers along mutually orthogonal directions, an object can be constrained to three mutually orthogonal planes, and thus be precisely located.
-Therefore, the first step to achieve the goal of this study is to identify grasp poses along three mutually orthogonal directions.
-We are going to introduce the grasp planning and grouping methods in the first subsection below.
-Once the grasp poses and groups are determined, search and motion planning techniques can be employed to generate regrasp actions and consequently the robot motion required to eliminate uncertainties.
-The second subsection below details the specific methods for search and motion planning.
-III-A Orthogonal Grasp Triples
+Both vision and visual-tactile information were used for learning.
+This study can be partially classified as research to reduce uncertainties in grasping poses during the regrasp planning process.
+However, our primary goal is not to achieve precise pose estimation for regrasping.
+Instead, we focus on actively leveraging the consecutive geometric constraints inherent in the regrasping process [25][26][27][28] to reduce uncertainty.
+From a broader perspective, our research also falls within the category of methods that utilize interactive perception [29] to estimate the pose of grasped objects.
+In addition to the grasping-based interaction approaches, recent literature on non-prehensile interaction has been receiving growing attention.
+For example, Jankowski et al.
+[30] discussed the stochastic contact dynamics of pushing and proposed a pushing planning method that minimizes uncertainty considering informed priors.
+Dong et al.
+[31], on the other hand, modeled the contact dynamics of pushing as an energy model and used constraints on the energy margin to plan robust pushing.
+III Regrasp Planning Using Orthogonal Triplets
 […]
@@
 Each color in the subfigure represents one group.
-Subsequently, based on the results of the grasp planning, we proceed by separating them into groups.
-Since our goal is to identify three mutually orthogonal grasp poses, we set our grouping criterion to be based on the opening and closing directions of the gripper fingers.
-Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 1 Note 1 Note 1In this context, the direction of $\boldsymbol{v}_{i}$ is not considered.
-That is, $\boldsymbol{v}_{i}$ assumed identical to $-\boldsymbol{v}_{i}$..
-Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …The candidate grasp poses in Fig.
-2(b) can be right classified into three mutually orthogonal groups.
+Subsequently, based on the results of the grasp planning, we separate the grasps into groups.
+Since our goal is to identify three mutually orthogonal grasp poses, the grouping criterion is defined by the opening and closing directions of the gripper fingers.
+Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 2 Note 2 Note 2In this context, the direction of $\boldsymbol{v}_{i}$ is not considered..
+Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, ….
+The candidate grasp poses in Fig.
+2(b) can be correctly classified into three mutually orthogonal groups.
 They are visualized in Fig.
 2(c) using red, green, and blue colors.
-After obtaining the grasp groups, we perform triple combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint.
-To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triple combination.
+After obtaining the grasp groups, we perform triplet combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint.
+To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triplet combination.
 $\mathtt{score}=|\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{j}|+|\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{k}|+|\boldsymbol{v}_{j}\cdot\boldsymbol{v}_{k}|$ (1)
-The $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ in the above equation represent the opening and closing directions of three different groups in a triple.
-We sort the scores of all triples and prioritize them with a score closest to 0 for regrasp planning and uncertainty elimination.
+The $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ in the above equation represent the opening and closing directions of three different groups in a triplet.
+We sort the scores of all triplets and prioritize them with a score closest to 0 for regrasp planning and uncertainty reduction.
 The closer the score value is to 00, the more orthogonal the three grasps from the groups are to each other.
-By performing successive regrasp using the prioritized triples, object uncertainties can be effectively minimized.
+By performing successive regrasp using the prioritized triplets, object uncertainties can be effectively minimized.
+Although we consistently refer to “orthogonal” grasps or groups, it is important to note that strict 90-degree orthogonality is not required to determine the object’s pose.
 The example shown in Fig.
 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero.
-In more general cases, there may be multiple candidate groups and triple combinations.
-The triples may have non-zero scores and regrasp should be performed by prioritizing the ones with smaller scores.
-III-B Regrasp Planning
-For the sorted triples, motion planning is sequentially performed on each of them, and the first successful solution is output as a valid candidate.
-Note that each triple comprises three groups of grasp poses.
+In more general cases, there may be multiple candidate groups and triplet combinations.
+The triplets may have non-zero scores, However, as long as the grasp direction vectors $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ form a non-singular matrix when arranged as columns, the pose can be estimated.
+This point will be clearer in Section V when we introduce the pose estimation algorithm.
+Here, the reason we sort triplet grasp combinations by their score is not for enforcing strict orthogonality, but rather to account for potential changes of grasp directions caused by subsequent admittance control.
+Prioritizing grasp triplets with higher orthogonality helps minimize the impact of grasp direction deviations.
+We will discuss this aspect in more detail in the admittance control section.
+After identifying orthogonal groups, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups.
+We use an incremental method to explore combinations of grasp triplets from the orthogonal groups and output a successful motion once found.
+Each triplet comprises three groups of grasp poses.
 During planning, one pose from each group needs to be selected to form a combination of three grasps for computation.
@@
 The grasps in each group are sequentially assigned to the two robotic arms for planning.
-In this way, we incrementally explore the combinations of the grasp triples, and output successful motion once found.
+In this way, we incrementally explore the combinations of the grasp triplets and output successful motion once found.
 The detailed algorithmic flow is illustrated in Fig.
 3.
-Note that the process illustrated in the figure is merely an incremental search.
-If a large amount of data can be accumulated using the proposed method in the future, we strongly recommend exploring learning-based methods to improve efficiency.
-Figure 3: Regrasp planning used in this study is a simple incremental search method, which sequentially traverses grouped grasp triples until a valid solution is found.
-Since the regrasp planning algorithm itself is not the primary focus of this paper, further optimization aspects such as acceleration or grasp selection refinement are not explored.
+Figure 3: Regrasp planning used in this study is a simple incremental search method, which sequentially traverses grouped grasp triplets until a valid solution is found.
+It should be noted that our grasping and planning methods are probabilistic, and their success rates inherently depend on the granularity, number, and diversity of the grasp samples.
+At the same time, increasing the number of grasps inevitably impacts planning time.
+Identifying the optimal trade-off between these factors is an important challenge that could be addressed using predictive approaches.
+If a large amount of data can be accumulated using the proposed method in the future, we are interested in exploring probabilistic modeling methods to learn grasp triplets and thus improve planning efficiency.
+However, this is beyond the scope of the present paper and is therefore not elaborated here.
 IV Conforming to Object Pose Uncertainty
@@
 Consider the scenario shown in Fig.
-4(a), where the left hand has already moved the object from its initial pose to the handover pose, and the right hand is approaching to perform the regrasp.
-The target pose of the right hand is planned based on an assumed pose in the simulation.
+4(a), where the giving hand has already moved the object from its initial pose to the handover pose, and the receiving hand is approaching to perform the regrasp.
+The target pose of the giving hand is planned based on an assumed pose in the simulation.
 However, due to uncertainties in the real world, the object’s assumed pose in the simulation deviates from its actual pose, as illustrated by the orange dashed frames and red solid frames in Fig.
 4(a).
-Consequently, the right hand’s grasping pose becomes misaligned with the actual object pose.
+Consequently, the receiving hand’s grasping pose becomes misaligned with the actual object pose.
 Applying position control for grasping in this scenario may lead to undesirable outcomes.
@@
 Therefore, rigidly executing the planned grasping motion solely with position control is not advisable.
-Figure 4: (a) The right hand approaches to grasp the object.
-Due to uncertainties, the object is not positioned exactly at the center of the two fingers as assumed but is instead slightly shifted upward.
+Figure 4: (a) The receiving hand approaches to grasp the object.
+Due to uncertainties, the object is not positioned exactly at the center of the two fingers, as assumed, but is instead slightly shifted upward.
 (b) At this point, if position control is used, the upper-side finger will make contact with the object first, exerting excessive force on it and causing deformation.
 (c) Comparison of object deformation before and after grasping.
-To address this issue, this study employs admittance control during robotic grasping and enables the manipulator’s pose to conform to the object’s pose.
-The basic concept of admittance control is to establish a mass-damper-spring system’s motion equation based on the external force exerted on the robot and the desired target force .
-The robot then follows the solution of this equation to adjust its motion accordingly, thereby accommodating external forces.
-For a desired robot position $x_{d}$, the admittance control equation is
-$M(\ddot{\mathit{x}}_{p}-\ddot{\mathit{x}}_{d})+B(\dot{\mathit{x}}_{p}-\dot{\mathit{x}}_{d})+K(\mathit{x}_{p}-\mathit{x}_{d})=\mathit{f}+\mathit{f}_{d}.$ (2)
-Here, $x_{p}$, $\dot{x_{p}}$, and $\ddot{x_{p}}$ are respectively the measured current robot end-effector position, velocity, and acceleration.
-$M$, $B$, $K$ denote the virtual inertia, damping, and stiffness parameters.
-$f$ is the force applied to the system, while $f_{d}$ represents the target force exerted by the robot after applying admittance control.
-In our specific grasping scenario, the target force is set to zero, as our objective is to ensure that the robot does not exert any external force on the object beyond the necessary grasping force, thereby preventing unintended movement, deformation, or damage.
-During the grasping process, when one of the robotic fingers first makes contact with the object, it receives a reaction force from the object, which corresponds to the external force $f$ in equation (2).
-Based on the robot’s current $x_{p}$, $\dot{x_{p}}$, $\ddot{x_{p}}$, as well as the force $f$ measured using a force/torque sensor installed at the robot manipulator’s wrist, we can compute the desired position $x_{d}$, velocity $\dot{x_{d}}$, and acceleration $\ddot{x_{d}}$ using the equation.
-These computed values are then used to control the robot, enabling it to conform to the object’s uncertain pose during the grasping process.
+To address this issue, this study employs admittance control during robotic grasping and enables the receiving hand’s grasping pose to conform to the uncertain object pose.
+Admittance control establishes a mass-damper-spring system’s motion equation based on the reaction force exerted on the receiving hand and the desired target force [6].
+The receiving hand follows the solution of the equation to adjust its motion accordingly, thereby accommodating reaction force.
+For a desired receiving hand position $\boldsymbol{x}_{d}$, the admittance control equation is
+$\boldsymbol{M}(\ddot{\boldsymbol{x}}_{p}-\ddot{\boldsymbol{x}}_{d})+B(\dot{\boldsymbol{x}}_{p}-\dot{\boldsymbol{x}}_{d})+K(\boldsymbol{x}_{p}-\boldsymbol{x}_{d})=\boldsymbol{f}+\boldsymbol{f}_{d}.$ (2)
+Here, $\boldsymbol{x}_{p}$, $\dot{\boldsymbol{x}_{p}}$, and $\ddot{\boldsymbol{x}_{p}}$ are respectively the measured current end-effector position, velocity, and acceleration of the receiving hand.
+$\boldsymbol{M}$, $\boldsymbol{B}$, and $\boldsymbol{K}$ denote the virtual inertia, damping, and stiffness parameters.
+$\boldsymbol{f}$ is the reaction force applied to the receiving hand when grasping.
+It can be measured using the force/torque sensor at the wrist of the receiving manipulator.
+$\boldsymbol{f}_{d}$ represents the target exerting force of the receiving hand after applying admittance control.
+It is essentially zero, as our goal is to ensure that the receiving hand does not exert any external force on the object beyond the necessary grasping force, thereby preventing unintended movement, deformation, or damage.
+Given the robot’s current $\boldsymbol{x}_{p}$, $\dot{\boldsymbol{x}_{p}}$, $\ddot{\boldsymbol{x}_{p}}$, $\boldsymbol{f}$, the desired trajectory $\boldsymbol{x}_{d}$, $\dot{\boldsymbol{x}}_{d}$, and $\ddot{\boldsymbol{x}}_{d}$ can be computed from the admittance control equation.
+These computed values are then fed to the admittance control law to drive the receiving hand to conform to the object’s uncertain pose during grasping.
 Fig.
@@
 In Fig.
-5(a), the upper-side finger first makes contact with the object and experiences a reaction force $f$.
+5(a), the upper-side finger first makes contact with the object and experiences a reaction force $\boldsymbol{f}$.
 The admittance control then adjusts the robotic hand by moving it in a direction that balances this force, as shown in Fig.
@@
 The adjustment continues until both fingers are in contact with the object and their grasping forces reach equilibrium.
-At this point, $f$ disappears and the grasping forces become internal forces within the whole gripper-object system.
+At this point, $\boldsymbol{f}$ disappears, and the grasping forces become internal forces within the whole gripper-object system.
 The final grasping result is depicted in Fig.
@@
 Figure 5: Admittance control helps avoid deformation or damage.
-(a) When the upper-side finger makes contact with the object, a reaction force $f$ is detected.
-(b) Admittance control adjusts the robot hand by moving it in a direction that balances $f$.
-(c) After admittance control, the object is firmly grasped, and no deformation occurs in its shape.
+(a) A reaction force $\boldsymbol{f}$ is detected when the upper-side finger contacts the object.
+(b) Admittance control adjusts the receiving hand by moving it to balance $\boldsymbol{f}$.
+(c) After admittance control, the object is firmly grasped without change.
+Note that the above equation describes only the position-based admittance control.
+A similar formulation also applies to rotational admittance control.
+However, the admittance controller may redirect the grasp toward an adjacent surface due to rotational errors.
+In extreme cases, this can result in parallel regrasps, making the object’s pose indeterminate.
+As a safeguard against such situations, we suggest prioritizing near-orthogonal grasps following equation (1).
+The design minimizes the likelihood that the three grasps lie on the same plane, even when one of the grasp directions is redirected due to unexpected admittance behavior.
 V Pose Estimation Using Conformed Grasp Poses
-Next, we present the method for estimating the object’s pose.
-In the planning simulation, we used an assumed object pose.
-This pose serves as the basis for all grasping and planning processes.
-Since there is no uncertainty in the simulation, the final object pose remains deterministic.
-Here, we denote the object’s final pose in the planning simulation as $\boldsymbol{p}_{o_{s}}$ and $\boldsymbol{R}_{o_{s}}$, and denote the three consecutive grasping poses of the robot in the simulation as $\boldsymbol{g}_{s_{1}}$, $\boldsymbol{g}_{s_{2}}$, and $\boldsymbol{g}_{s_{3}}$.
-Each $\boldsymbol{g}_{s_{i}}$ consists of ($\boldsymbol{p}_{s_{i}}$, $\boldsymbol{R}_{s_{i}}$, $d_{s_{i}}$), where $\boldsymbol{p}_{s_{i}}$ and $\boldsymbol{R}_{s_{i}}$ represent the position and rotation of the grasping center point, and $d_{s_{i}}$ denotes the jaw opening width.
-On the other hand, by executing the planned motion in the real world, we obtain three consecutive grasping poses of the bimanual robot in the real world as $\boldsymbol{g}_{r_{1}}$, $\boldsymbol{g}_{r_{2}}$, and $\boldsymbol{g}_{r_{3}}$.
-Each $\boldsymbol{g}_{r_{i}}$ similarly consists of ($\boldsymbol{p}_{r_{i}}$, $\boldsymbol{R}_{r_{i}}$, $d_{r_{i}}$).
-The first grasping pose $\boldsymbol{g}_{r_{1}}$ aligns exactly with the planned result, while the second grasping pose $\boldsymbol{g}_{r_{2}}$ and the third grasping pose, $\boldsymbol{g}_{r_{3}}$, are adjusted through admittance control to conform to the object’s actual pose.
-They are different from $\boldsymbol{g}_{s_{2}}$ and $\boldsymbol{g}_{s_{3}}$.
-Our objective is to estimate the real-world object pose $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ based on these known values.
-To address this issue, let us first consider the simulation scenario.
-Since our three grasping poses involve planar contacts and are mutually orthogonal, the object constrained by them should have a unique pose.
-The pose can be computed as follows.
-For each grasp, we can obtain a corresponding opening and closing vector $\boldsymbol{v}_{i}=\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$.
-Here, $\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$ denotes the $\boldsymbol{y}$ axis (2nd column) of the rotation matrix.
-We assume the robot hand’s opening and closing direction is aligned with its local $\boldsymbol{y}$ axis.
-Since the three grasping poses are mutually orthogonal, they uniquely determine the following rotation matrix
-$\boldsymbol{R}_{g_{s}}=\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})&\boldsymbol{R}_{s_{2}}(\boldsymbol{y})&\boldsymbol{R}_{s_{3}}(\boldsymbol{y})\end{bmatrix}.$ (3)
-Meanwhile, for each grasp, its two parallel fingers define a midplane.
-Given that the three grasps are mutually orthogonal, their corresponding midplanes intersect at a unique intersection point.
-This intersection represents a unique position constraint imposed by the grasping configuration.
-It can be computed using
-$\boldsymbol{p}_{g_{s}}=(\boldsymbol{R}_{g_{s}}^{\top})^{-1}\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{1}}}\\ \boldsymbol{R}_{s_{2}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{2}}}\\ \boldsymbol{R}_{s_{3}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{3}}}\end{bmatrix}$ (4)
-The ($\boldsymbol{p}_{g_{s}}$, $\boldsymbol{R}_{g_{s}}$) obtained from the above two equations are the unique pose determined by the three grasping poses in the planning simulation.
-Note that it is not the object’s final pose ($\boldsymbol{p}_{o_{s}}$, $\boldsymbol{R}_{o_{s}}$).
-However, since both are uniquely determined.
-There is a unique offset between them written as
-$\boldsymbol{p}_{\text{diff}}=\boldsymbol{p}_{o_{s}}-\boldsymbol{p}_{g_{s}},~{}\boldsymbol{R}_{\text{diff}}=\boldsymbol{R}_{g_{s}}^{\top}\boldsymbol{R}_{o_{s}}$ (5)
-Fig. 6 illustrates the three planes formed by three grasping poses and their intersection point.
-Figure 6: Three orthogonal grasping poses with planar contacts determine three planes that intersect at a unique point.
-Thus, a unique position $\boldsymbol{p}_{g_{s}}$ is determined.
-Meanwhile, the opening-closing directions of the three gasping poses help constrain a unique rotation $\boldsymbol{R}_{g_{s}}$
-In the real world, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ are uniquely constrained by the grasping poses and have the same offset from the pose determined by the grasping poses. That is
-$\boldsymbol{p}_{o_{r}}=\boldsymbol{p}_{g_{r}}+\boldsymbol{p}_{\text{diff}},~{}\boldsymbol{R}_{o_{r}}=\boldsymbol{R}_{g_{r}}\boldsymbol{R}_{\text{diff}}.$ (6)
-Thus, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ can be computed.
-Here, $\boldsymbol{p}_{g_{r}}$ and $\boldsymbol{R}_{g_{r}}$ can be obtained similarly to equations (4) and (3) using ($\boldsymbol{p}_{r_{1}}$, $\boldsymbol{R}_{r_{1}}$), ($\boldsymbol{p}_{r_{2}}$, $\boldsymbol{R}_{r_{2}}$), and ($\boldsymbol{p}_{r_{3}}$, $\boldsymbol{R}_{r_{3}}$).
+V-A Error Parameterization
+We denote the pose of the object and first grasp in the world frame of the simulation as $({}^{w}\mathbf{p}_{o}^{\text{sim}},{}^{w}\mathbf{R}_{o}^{\text{sim}})$, and $({}^{w}\mathbf{p}_{g_{1}}^{\text{sim}},{}^{w}\mathbf{R}_{g_{1}}^{\text{sim}})$, respectively.
+The relative pose of the object with respect to the first grasp can be computed as
+{${}^{g_{1}}\mathbf{R}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{1}}^{\text{sim}})^{\top}\,{}^{w}\mathbf{R}_{o}^{\text{sim}}\\ {}^{g_{1}}\mathbf{p}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{1}}^{\text{sim}})^{\top}({}^{w}\mathbf{p}_{o}^{\text{sim}}-{}^{w}\mathbf{p}_{g_{1}}^{\text{sim}})\end{aligned}\right..$ (3)
+Meanwhile, we denote the first grasp’s actual pose after impedance control as $({}^{w}\mathbf{p}_{g_{1}}^{\text{real}},{}^{w}\mathbf{R}_{g_{1}}^{\text{real}})$.
+Without considering uncertainty, the pose of the object in the world frame of the real world can be computed as
+{${}^{w}\mathbf{R}_{o}^{\text{ideal}}({g_{1}})&={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}\cdot{}^{g_{1}}\mathbf{R}_{o}^{\text{sim}}\\ {}^{w}\mathbf{p}_{o}^{\text{ideal}}({g_{1}})&={}^{w}\mathbf{p}_{g_{1}}^{\text{real}}+{}^{w}\mathbf{R}_{g_{1}}^{\text{real}}\cdot{}^{g_{1}}\mathbf{p}_{o}^{\text{sim}}\end{aligned}\right.,$ (4)
+The subscript $(g_{1})$ indicates the computations are based on the first grasp. The superscript “ideal” indicates the results are an estimation without considering uncertainty.
+To account for uncertainty introduced during the grasping process, we model the object pose error as comprising a rotational component $\mathbf{R}_{o}^{\text{err}}({g_{1}})$ and a translational component $\mathbf{p}_{o}^{\text{err}}({g_{1}})$.
+Similarly, $(g_{1})$ indicates that the errors are associated with the first grasp.
+Due to the geometric constraints between the gripper and the object surface, $\mathbf{R}_{o}^{\text{err}}({g_{1}})$ is constrained to a rotation about the second column of the real-world grasp frame, ${}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}[:,1]$, and $\mathbf{p}_{o}^{\text{err}}({g_{1}})$ is restricted to the plane spanned by the first and third columns ${}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}}={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}[:,0]$ and ${}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}}={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}[:,2]$.
+The three arrows in Fig.
+6(a.ii) show these columns.
+The errors can thus be parameterized as
+$\mathbf{R}_{o}^{\text{err}}({g_{1}})&=\exp\left(\theta\,[{}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}]_{\times}\right)\\ \mathbf{p}_{o}^{\text{err}}({g_{1}})&=\delta_{g_{1,1}}\,{}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}}+\delta_{g_{1,3}}\,{}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}}\end{aligned}\right..$ (5)
+Accordingly, the true pose of the object in the real world is
+{${}^{w}\mathbf{R}_{o}^{\text{real}}&={}^{w}\mathbf{R}_{o}^{\text{ideal}}(g_{1})\cdot\mathbf{R}_{o}^{\text{err}}({g_{1}})\\ {}^{w}\mathbf{p}_{o}^{\text{real}}&={}^{w}\mathbf{p}_{o}^{\text{ideal}}(g_{1})+\mathbf{p}_{o}^{\text{err}}({g_{1}})\end{aligned}\right..$ (6)
+[…]
+V-B Rotation Error Identification
+We can use the second grasp to identify $\mathbf{R}_{o}^{\text{err}}({g_{1}})$.
+Let $({}^{w}\mathbf{p}_{g_{2}}^{\text{sim}},{}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})$ denote the simulated pose of the second grasp (equally the grasp pose before admittance control), and let $({}^{w}\mathbf{p}_{g_{2}}^{\text{real}},{}^{w}\mathbf{R}_{g_{2}}^{\text{real}})$ denote its real-world pose after admittance control.
+The rotation difference from the simulation to the real execution is $\mathbf{R}_{g_{2}}^{\text{diff}}=({}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})^{\top}\cdot{}^{w}\mathbf{R}_{g_{2}}^{\text{real}}$, which captures the rotational deviation introduced by the admittance control.
+Since $\mathbf{R}_{o}^{\text{err}}({g_{1}})$ is constrained to only occur around $\mathbf{r}_{g_{1,2}}^{\text{real}}$, we extract the component of $\mathbf{R}_{g_{2}}^{\text{diff}}$ along $\mathbf{r}_{g_{1,2}}^{\text{real}}$ to get $\theta$:
+$\theta=\text{axis\_angle}(\mathbf{R}_{g_{2}}^{\text{err}},{}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}),$ (7)
+and consequently $\mathbf{R}_{o}^{\text{err}}({g_{1}})$. The object’s true rotation in the real world can be uniquely identified as
+${}^{w}\mathbf{R}_{o}^{\text{real}}={}^{w}\mathbf{R}_{o}^{\text{ideal}}({g_{1}})\cdot\mathbf{R}_{o}^{\text{err}}({g_{1}})={}^{w}\mathbf{R}_{o}^{\text{ideal}}({g_{1}})\cdot\exp\left(\theta\,[{}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}]_{\times}\right).$ (8)
+The rotation of the purple object in Fig. 6(a.iii) illustrates an identified ${}^{w}\mathbf{R}_{o}^{\text{real}}$.
+V-C Translation Error Identification
+Next, we address $\mathbf{p}_{o}^{\text{err}}({g_{1}})$. In simulation, the pose of the object relative to the second grasp is
+{${}^{g_{2}}\mathbf{R}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})^{\top}\,{}^{w}\mathbf{R}_{o}^{\text{sim}}\\ {}^{g_{2}}\mathbf{p}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})^{\top}({}^{w}\mathbf{p}_{o}^{\text{sim}}-{}^{w}\mathbf{p}_{g_{2}}^{\text{sim}})\end{aligned}\right..$ (9)
+Let $({}^{w}\mathbf{p}_{g_{2}}^{\text{real}},{}^{w}\mathbf{R}_{g_{2}}^{\text{real}})$ be the second grasp’s real-world pose.
+If there is no translational uncertainty, the position of the object in the real world can be computed using the second grasp as
+${}^{w}\mathbf{p}_{o}^{\text{ideal}}({g_{2}})={}^{w}\mathbf{p}_{g_{2}}^{\text{real}}+{}^{w}\mathbf{R}_{g_{2}}^{\text{real}}\,{}^{g_{2}}\mathbf{p}_{o}^{\prime}.$ (10)
+Since ${}^{w}\mathbf{R}_{o}^{\text{real}}$ has already been identified, the corrected relative position ${}^{g_{2}}\mathbf{p}_{o}^{\prime}$ instead of ${}^{g_{2}}\mathbf{p}_{o}^{\text{sim}}$ is used in the equation.
+${}^{g_{2}}\mathbf{p}_{o}^{\prime}$ is computed as
+${}^{g_{2}}\mathbf{p}_{o}^{\prime}={}^{g_{2}}\mathbf{R}_{o}^{\text{real}}\left({}^{g_{2}}\mathbf{R}_{o}^{\text{sim}}\right)^{\top}{}^{g_{2}}\mathbf{p}_{o}^{\text{sim}},$ (11)
+where
+${}^{g_{2}}\mathbf{R}_{o}^{\text{real}}=({}^{w}\mathbf{R}_{g_{2}}^{\text{real}})^{\top}\,{}^{w}\mathbf{R}_{o}^{\text{real}}.$ (12)
+The translational uncertainty caused by the second grasp lies within the plane spanned by the first and third columns of ${}^{w}\mathbf{R}_{g_{2}}^{\text{real}}$ and can be parameterized as
+$\mathbf{p}^{\text{err}}({g_{2}})=\delta_{2,1}\,{}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}}+\delta_{2,3}\,{}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}},$ (13)
+where ${}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}}={}^{w}\mathbf{R}_{g_{2}}^{\text{real}}[:,0]$ and ${}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}}={}^{w}\mathbf{R}_{g_{2}}^{\text{real}}[:,2]$.
+Fig.
+6(a.iii) illustrates them.
+The true position of the object can be expressed using $\mathbf{p}^{\text{err}}({g_{2}})$ as
+${}^{w}\mathbf{p}_{o}^{\text{real}}={}^{w}\mathbf{p}_{o}^{\text{ideal}}(g_{2})+\mathbf{p}_{o}^{\text{err}}({g_{2}}).$ (14)
+It is important to note that the translational errors associated with the first and second grasps are respectively constrained to planes determined by $({}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}},{}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}})$ and $({}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}},{}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}})$.
+Their normal vectors are $\mathbf{n}_{g_{1}}={}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}}\times{}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}}$ and $\mathbf{n}_{g_{2}}={}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}}\times{}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}}$.
+Since the translational error must simultaneously satisfy the constraints imposed by both planes, it is restricted to lie along their intersection line.
+The direction of this line can be computed using
+$\mathbf{d}_{\text{allowed}}=\frac{\mathbf{n}_{g_{1}}\times\mathbf{n}_{g_{2}}}{\|\mathbf{n}_{g_{1}}\times\mathbf{n}_{g_{2}}\|}.$ (15)
+Thus, the translational error can essentially be parameterized using a single scalar value as $\epsilon\,\mathbf{d}_{\text{allowed}}$, and the true real-world object position simplifies to
+${}^{w}\mathbf{p}_{o}^{\text{real}}={}^{w}\mathbf{p}_{o}^{\text{ideal}}({g_{2}})+\epsilon\,\mathbf{d}_{\text{allowed}}.$ (16)
+The scalar value $\epsilon$ can be easily determined from the displacement of the third grasp after admittance control.
+Let ${}^{w}\mathbf{p}_{g_{3}}^{\text{sim}}$ and ${}^{w}\mathbf{p}_{g_{3}}^{\text{real}}$ denote the simulated position of the third grasp (position before admittance control) and the real-world position after admittance control, respectively.
+The displacement is computed as ${}^{w}\mathbf{p}_{g_{3}}^{\text{real}}-{}^{w}\mathbf{p}_{g_{3}}^{\text{sim}}$.
+The scalar $\epsilon$ can be obtained using the following projection
+$\epsilon=\mathbf{d}_{\text{allowed}}^{\top}({}^{w}\mathbf{p}_{g_{3}}^{\text{real}}-{}^{w}\mathbf{p}_{g_{3}}^{\text{sim}})),$ (17)
+as indicated within the circle in Fig. 6(a.iv). Substituting the obtained $\epsilon$ into equation (16) yields the object’s true position.
+In addition to the L-shaped object with a square cross-section discussed thus far, Fig.
+6(b) presents an L-shaped object with a diamond cross-section.
+Although this object does not contain any pair of strictly orthogonal grasping directions, it is still possible to select three non-parallel grasps and estimate the final object pose using the method described in this section.
+Fig.
+6(b.i) shows the uncertainty after the first grasp, while Fig.
+6(b.ii) illustrates the orientation determined after the second grasp, along with the direction of the remaining translational error.
+A detailed 3D visualization of both the L-shaped objects can be found in the supplementary video.
 VI Experiments and Analysis
-This section presents the real-world validation of the proposed method.
-Our experimental platform consists of two UR3e robots from Universal Robots.
-Their installation setup is shown in Fig.
+Our experimental platform consists of two UR3e robots.
+Their setup is shown in Fig.
 7.
-The UR3e series robots are equipped with high-precision force/torque sensors at their end effectors, which facilitate the implementation of admittance control.
-Each robot is installed with a Robotiq Hand-E gripper from Robotiq.
+Each robot is installed with a Robotiq Hand-E gripper.
 The opening range of the gripper is 0$\sim$50 mm.
-The regrasp planning is performed using the WRS simulation environment Note 2 Note 2 Note 2https://github.com/wanweiwei07/wrs.
+The regrasp planning is performed using the WRS simulation environment.
+The coordinate system of the platform is defined at the back of the robot, with its origin at the intersection of the central plane between the two robot arms and the ground.
+The $\boldsymbol{x}$-axis points toward the monitor in the figure.
+The $\boldsymbol{y}$-axis points towards the first arm.
+The $\boldsymbol{z}$-axis points upward.
+For comparison, we attached optical markers to both the objects and the robots, and tracked the poses of the resulting rigid bodies using three tracking cameras.
+The poses estimated by the proposed method were compared with those captured by the optical tracking system to evaluate the method’s performance.
+The tracking system provides a detection accuracy of $\pm$0.5 mm.
+[…]
 Fig.
-6 in the previous section provided a view of the simulation environment.
-The coordinate system of our experimental platform is defined with its origin at the intersection of the central plane between the two robot arms and the rear edge of the table, as illustrated by the colored frame in Fig.
-7.
-The $\boldsymbol{x}$-axis points forward, the $\boldsymbol{y}$-axis points towards the left arm, and the $\boldsymbol{z}$-axis aligns with the normal direction of the tabletop surface.
-[…]
-To ensure clarity in comparison and analysis, we used the simple L-shaped object shown in Fig.
-7 and assumed a single initial pose in the planning simulation.
-The object has a long edge of $125$ mm and a short edge of $100$ mm, with a square cross-section of $25$ mm $\times$ $25$ mm.
-The surfaces of the L-shaped object are mutually perpendicular.
-They could align with the three axes of the coordinate system and facilitate analysis.
-The initial pose of the object in the simulation was set to $(700,300,1.25)$ mm.
-Its rotation was set to be aligned with the coordinate frame of the robot platform (recall the colored frame shown in Fig.
-7).
-The long edge aligns with the $\boldsymbol{y}$ axis.
-The short edge aligns with the $\boldsymbol{z}$ axis.
-Similarly to the initial pose, we set the regrasp pose of the object in the simulation to be at a unique fixed location $(680,50.0,120.0)$ mm, with the rotation identical to the initial one to facilitate analysis.
-The location is in the middle of the two arms, making it easy to plan effective bimanual regrasp motions.
-The regrasp motion planned based on the parameter settings and the proposed planning methods led to orthogonal triples where the first selected grasp is along the $\boldsymbol{y}$ axis, the second grasp is along the $\boldsymbol{z}$ axis, and the third grasp is along the $x$ axis.
-Together, the three grasps uniquely constrain the object’s pose.
-In the real world, we manually introduced uncertainties in multiple directions to the object for experimental evaluation.
-These uncertainties are illustrated in Fig.
-8.
-Fig.
-8(a) represents initial object poses obtained by introducing uncertainties in the horizontal directions (back, front, left, and right) relative to a center pose shown in Fig.
-8(a.v).
-The center pose closely corresponds to the initial pose of the object used in the simulation.
-Note that the center pose is not exactly the initial pose in the simulation.
-The front, back, left, and right movements are approximate ±$\pm 12.5$ mm and are also not exact.
-We are not enforcing precise placement of the object when preparing these poses.
-The reason is that preparing exact uncertainty in the real world can be challenging.
-Finger contact during the grasping process naturally causes additional errors.
-Specifying exact uncertainty values and evaluating absolute accuracy would not provide meaningful insights.
-We analyze repeatability errors rather than absolute errors to evaluate the effectiveness of our method in eliminating uncertainties.
-In addition to the horizontal uncertainties, we also prepared initial poses with uncertainties in the upward, upper-front, and upper-back directions.
-These poses are shown in Fig.
-8(b).
-The L-shape object is elevated using a white block while being shifted back and forward with the same approximate ±$\pm 12.5$ mm offset when preparing these uncertain poses.
-Figure 8: (a) $9$ uncertain poses in the horizontal directions.
-(a.i) BL – Back Left.
-(a.ii) BC – Back Center.
-(a.iii) BR – Back Right.
-(a.iv) L – Left.
-(a.v) C – Center.
-(a.vi) R – Right.
-(a.vii) FL – Front Left.
-(a.viii) FC – Front Center.
-(a.ix) FR – Front Right.
-(b) $3$ uncertain poses in upward directions.
-(b.i) DC – Down Center (for reference).
-(b.ii) UC – Upper Center.
-(b.iii) Upper Back.
-(b.iv) Upper Front.
-Using the planned robot motions, we performed regrasping and applied admittance control following the proposed method for each initial object pose in Figs.
-8(a) and (b), and estimated their final object poses using equation 6.
-The estimation results are summarized in the “Proposed method” columns of Table I.
-The results corresponding to (a) and (b) are separated by a double solid line.
-From the results, we can observe that the estimation performs reasonably well.
-In the case of the uncertain object placement shown in Figs.
-8(a), the estimated $\boldsymbol{p}_{o_{r}}$ vector has a mean and variance of $(9.82,0.002)$ mm and $(898.33,0.019)$ mm in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions respectively.
-The low variance aligns well with our intuition – For the uncertain poses listed in Fig.
-8(a), the object is placed on the table.
-Its position in the $\boldsymbol{z}$-direction remains the same.
-Meanwhile, the object is always grasped by fingers closing from the $\boldsymbol{y}$-direction.
-Its uncertainty in the $\boldsymbol{y}$-direction is constrained first grasp’s two fingers.
-Therefore, the final estimated poses should be consistent (have a low variance) in these two directions.
-On the other hand, the first values of the estimated $\boldsymbol{p}_{o_{r}}$ can be clearly categorized into three distinct groups.
-Among them, BC and BR form one group, with a mean and variance of $(687.05,2.1)$ mm.
-L, C, and R constitute another group, with a mean and variance of $(699.3,1.4)$ mm.
-Finally, FL and FC belong to the third group, with a mean and variance of $(709.45,2.1)$ mm.
-This result is also consistent with our initial placement.
-The object displacement in the $\boldsymbol{x}$ direction can be clearly distinguished, and the repeatability is satisfying.
-The variance of each distinct group falls within the range of the manual placement uncertainty.
-Similar analysis and conclusions can be drawn from the results for the uncertain object placements shown in Figs.
-8(b).
-They are in the UB, UC, and UF rows.
-The estimated $\boldsymbol{p}_{o_{r}}$ vectors have nearly the same values in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions and clearly distinct values in the $\boldsymbol{x}$ direction.
-For rotation, all estimated results are generally aligned with the coordinate axes.
-They have an average deviation of $0.59^{\circ}$ with $0.11^{\circ}$ variance from rotation matrix $\begin{bmatrix}\boldsymbol{x}&\boldsymbol{z}&-\boldsymbol{y}\end{bmatrix}$.
-This outcome is consistent with the fact that the object’s initial placement is always on a plane and that the opening-closing direction of the first grasp is along the $\boldsymbol{y}$ axis.
-Its repeatability is also satisfying considering the small variance.
-Note that the BL and FR rows in the table do not contain values.
-This is because the object’s pose uncertainty was too large, causing the robot to collide with the object during execution.
-The maximum tolerable uncertainty in our method is constrained by the gripper’s opening width.
-When the uncertainty exceeds half of the gripper’s opening width, collisions may occur and lead to execution failure.
-Additionally, the table does not include estimation results for DC pose.
-This pose is only provided to give readers a reference for comparison before the object is elevated, and thus, no estimation was performed for it.
-TABLE I: Pose estimation results
-Type Exec. Proposed method AR marker
-$\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$ $\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$
-BL Fail. - - - -
-BC Succ.
-$(685.6,49.8,118.2)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.01&0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(715.4,56.5,128.3)$ $\begin{bmatrix}1.00&-0.01&0.07\\ 0.07&-0.02&-1.00\\ 0.01&1.00&-0.02\end{bmatrix}$
-BR Succ.
-$(688.5,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.02&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(719.5,34.8,146.3)$ $\begin{bmatrix}1.00&-0.06&0.00\\ 0.00&-0.02&-1.00\\ 0.06&1.00&-0.02\end{bmatrix}$
-L Succ.
-$(699.5,49.9,118.5)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(740.2,54.1,131.5)$ $\begin{bmatrix}1.00&-0.03&0.05\\ 0.05&-0.01&-1.00\\ 0.03&1.00&-0.00\end{bmatrix}$
-C Succ.
-$(697.8,49.9,118.1)$ $\begin{bmatrix}1.00&-0.00&0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(726.9,36.8,144.5)$ $\begin{bmatrix}1.00&-0.07&-0.01\\ -0.01&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
-R Succ.
-$(700.7,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(726.3,55.6,145.8)$ $\begin{bmatrix}1.00&-0.07&0.07\\ 0.06&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
-FL Succ.
-$(710.9,49.8,118.4)$ $\begin{bmatrix}1.00&0.01&-0.00\\ 0.03&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(741.3,57.7,128.4)$ $\begin{bmatrix}0.99&-0.01&0.10\\ 0.10&-0.01&-0.99\\ 0.01&1.00&-0.01\end{bmatrix}$
-FC Succ.
-$(708.0,49.8,118.5)$ $\begin{bmatrix}1.00&0.00&-0.00\\ 0.02&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(756.1,29.4,110.6)$ $\begin{bmatrix}1.00&-0.00&0.01\\ 0.01&-0.03&-1.00\\ 0.00&1.00&-0.03\end{bmatrix}$
-FR Fail. - - - -
-UB Succ.
-$(689.8,49.8,131.8)$ $\begin{bmatrix}1.00&0.01&0.00\\ -0.02&-0.01&-1.00\\ -0.03&1.00&0.00\end{bmatrix}$ $(717.5,38.3,157.9)$ $\begin{bmatrix}1.00&-0.06&-0.02\\ -0.02&-0.03&-1.00\\ 0.06&1.00&-0.03\end{bmatrix}$
-UC Succ.
-$(698.7,49.8,131.6)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(732.4,32.1,160.6)$ $\begin{bmatrix}1.00&-0.07&0.02\\ 0.02&-0.02&-1.00\\ 0.07&1.00&-0.02\end{bmatrix}$
-UF Succ.
-$(706.6,49.8,131.6)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ 0.01&-0.02&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(735.6,36.6,158.5)$ $\begin{bmatrix}1.00&-0.08&0.02\\ 0.01&-0.02&-1.00\\ 0.08&1.00&-0.02\end{bmatrix}$
-[…]
+8 shows the objects used in the experiments, their dimensions, and the randomly assigned placements in experimental trials.
+The objects are L-shaped items with square and diamond cross-sections, as described earlier.
+Their long arm measures 125 mm, the short arm 100 mm, and the distance between the parallel edges of the cross-section is 25 mm for both.
+The diamond-shaped cross-section has a 75 acute interior angle.
+Optical tracking markers were attached at pre-drilled holes on the objects.
+During grasp planning, we can automatically generate grasps and motions that avoid collisions with the markers.
+For each object, we randomly assigned three different initial placements, labeled L1$\sim$3 and TL1$\sim$3 in Fig.
+8, and tasked the robots with moving the objects to predefined target positions.
+The supplementary video accompanying this paper contains details of the grasp selection, motion planning, and robot actions.
+We recommend readers to refer to this video for an intuitive understanding of the experimental procedures.
+Figure 8: (a) L-shape object and its random initial placements. (b) TL-shape object and its random initial placements.
+Table I summarizes the results.
+The data section has two main column groups reporting the positional differences and rotational differences.
+Within each group, there are three subcolumns.
+The $g_{2}$ columns show the positional and rotational differences of the $g_{2}$ grasp before and after admittance control.
+The $g_{3}$ columns present the corresponding differences for the $g_{3}$ grasp.
+The “object” columns report the differences between the object pose estimated by the proposed method and the pose measured by the optical tracking system.
+The results for $g_{1}$ are not included, as it involves no admittance control.
+TABLE I: Experimental Results with Statistics
+Task Position Differences ($\Delta\mathbf{p}$, mm) Rotation Vector Differences ($\Delta\mathbf{w}$, degrees)
+$g_{2}$ $g_{3}$ object $g_{2}$ $g_{3}$ object
+L1 -0.40 -1.60 5.20 -8.90 -2.00 -0.80 -33.70 -61.80 32.20 -0.55 -0.91 -0.02 -0.34 -0.75 -0.91 -0.01 2.82 -1.39
+L2 -0.30 -1.70 5.20 -0.20 -2.00 -0.80 -33.60 -61.70 32.30 -0.51 -0.14 -0.03 -0.33 -0.25 -0.06 -0.12 1.83 -1.95
+L3 -0.40 -1.60 5.20 -9.00 -1.90 -0.40 -33.20 -61.50 31.70 -0.30 -0.13 -0.00 -0.20 -0.51 -1.11 -0.00 0.98 -1.07
+Mean -33.50 -61.67 32.07 -0.04 1.87 -1.47
+Std -00.25 -00.15 00.25 -0.06 0.95 -0.44
+TL1 -1.81 -0.99 0.38 -8.02 -8.05 -7.20 -59.35 -32.75 52.41 -0.64 -1.07 -1.00 -0.45 -0.36 -0.78 -0.01 2.05 -0.42
+TL2 -0.08 -0.18 1.31 -0.40 -1.60 5.20 -59.66 -34.42 51.76 -0.00 -0.00 -0.00 -0.30 -0.13 -0.00 -0.13 0.73 -1.09
+TL3 -0.07 -0.21 1.17 -1.71 -0.36 -1.47 -56.37 -36.34 52.26 -0.04 -0.01 -0.00 -1.25 -0.29 -0.56 -0.86 0.27 -2.11
+Mean -58.46 -34.50 52.14 -0.33 1.02 -1.21
+Std -01.86 -01.50 00.36 -0.47 0.38 -0.85
+From the table, we observe that both the $g_{2}$ and $g_{3}$ grasps exhibit substantial changes before and after admittance control, indicating that the robot actively adjusts the regrasp pose to estimate uncertainties for each task.
+However, the final estimation results are not satisfactory, showing positional differences of up to 61.67 mm and rotational differences of up to 1.86 degrees compared to the optical tracking system.
+Upon closer examination, we found that these errors are primarily systematic.
+For the L-shaped object, the standard deviations in $\Delta\mathbf{p}$ across tasks were [0.25, 0.15, 0.25] mm, all below 0.3 mm, and the deviations in $\Delta\mathbf{w}$ were [0.06, 0.95, 0.44] degrees, all below 1.0 degrees.
+Similarly, for the TL-shaped object, the deviations in $\Delta\mathbf{p}$ and $\Delta\mathbf{w}$ were [1.86, 1.50, 0.36] mm and [0.47, 0.38, 0.85] degrees, respectively, remaining below 2 mm and 0.9 degrees.
+These results suggest that while the overall estimation errors appear large, their internal consistency is high, indicating that the system’s inaccuracies stem largely from systematic biases.
+Likely sources include errors from the optical tracking system itself (due to the 3D-printed marker mounts for calibration and robot frame location), installation errors in the dual-arm robot, and absolute control errors of the robot.
+In addition, when comparing the deviations between the L-shaped and TL-shaped objects, we found that the inference results for the L-shaped object were better than those for the TL-shaped object.
+This outcome differs from our theoretical expectations.
+Theoretically, even if the three grasps are not perfectly orthogonal, the final inference accuracy should not degrade, and the L- and TL-shaped objects should exhibit similar deviations.
+We believe that the observed difference arises because the TL-shaped object involves rotations around axes unaligned with body frames, which place higher demands on the assembly precision of the dual-arm robot.
+The larger deviations reveal that the robot setup has greater misalignment in rotational components.
+In conclusion, although the results exhibit large absolute errors, the deviation levels are on the same order of magnitude as the optical tracking system.
+The results indicate that the proposed method has strong relative pose estimation performance.
 VII Conclusions and Future Work
-This paper proposes a method for eliminating object uncertainties by leveraging the grasping constraints of parallel grippers with flat finger pads and a bimanual regrasp planning method.
+This paper proposes a method for reducing object uncertainties by leveraging the grasping constraints of parallel grippers with flat finger pads and a bimanual regrasp planning method.
 We applied this method to estimate grasping errors for a roughly placed L-shaped object.
-The results demonstrate that the proposed approach effectively eliminates uncertainties and achieves high repeatability accuracy.
-Currently, this method is only applicable when all three grasps involve planar contact.
-For curved surface contacts, additional grasps may be required to fully constrain uncertainties.
-Further exploring the necessary grasp combinations will be an important future topic.
+The results demonstrate that the proposed approach can effectively reduce uncertainties and achieves high repeatability accuracy.
+The current method assumes planar contacts between flat fingertips and flat object surfaces.
+When the object surface is curved, the number of degrees of freedom constrained by a grasp will vary, and more grasps may be required to constrain uncertainties.
+Investigating the minimal and sufficient set of grasps under such conditions is an interesting future direction.
+Additionally, exploring learning-based methods to predict grasps and motion will be helpful to improve efficiency.

### v4

@@
 Our method is based on the concept that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact.
-Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state.
-Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to actively reduce uncertainties in the object pose.
+Three approximately orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state.
+Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three approximately orthogonal grasps of two robotic arms to actively reduce uncertainties in the object pose.
 We evaluated the proposed method on different initial object uncertainties and verified that it had good repeatability.
@@
 (b) Flat finger pads of a parallel gripper help constrain uncertainty to a plane.
-(c) Three orthogonal parallel grasps constrain three planes, leading to an exact pose.
+(c) Three aprpoximately orthogonal parallel grasps constrain three planes, leading to an exact pose.
 This paper proposes a bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to reduce uncertainties.
@@
 iii) The maximum uncertainty in the object pose is within the gripper’s opening range, ensuring valid contact during grasping..
-Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triplets to simultaneously reduce uncertainties in the object pose.
+Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting approximately orthogonal grasp triplets to simultaneously reduce uncertainties in the object pose.
 However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage.
-To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose.
-By integrating regrasp planning with admittance control, our method iteratively reduces uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
+To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors at the robots’ wrists to adjust the grasping pose and conform to the uncertain object pose.
+By integrating regrasp planning with admittance control, our method iteratively reduces uncertainties along three approximately orthogonal axes, thus achieving an accurate object pose estimation.
 In the experiments, we implemented the proposed method and evaluated its performance on different initial object uncertainties.
@@
 [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds.
-The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
+The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it falls under the first type of regrasp planning.
+It can be seen as a concrete realization of a flipped funnel [17], where the object, through interaction, progressively guides the three structured grasping poses, allowing its uncertain state to be estimated toward a determined pose.
 II-B Key Challenges and the Focus of this Work
@@
 (b) Candidate grasp poses planned using the antipodal method.
-(c) Separating the candidate grasp poses into three mutually orthogonal groups considering their opening and closing directions.
+(c) Separating the candidate grasp poses into three orthogonal groups considering their opening and closing directions.
 Each color in the subfigure represents one group.
 Subsequently, based on the results of the grasp planning, we separate the grasps into groups.
-Since our goal is to identify three mutually orthogonal grasp poses, the grouping criterion is defined by the opening and closing directions of the gripper fingers.
+Since our goal is to identify three approximately orthogonal grasp poses, the grouping criterion is defined by the opening and closing directions of the gripper fingers.
 Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 2 Note 2 Note 2In this context, the direction of $\boldsymbol{v}_{i}$ is not considered..
@@
 The candidate grasp poses in Fig.
-2(b) can be correctly classified into three mutually orthogonal groups.
+2(b) are classified into three strictly orthogonal groups.
 They are visualized in Fig.
-2(c) using red, green, and blue colors.
-After obtaining the grasp groups, we perform triplet combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint.
+2(c) using red, green, and blue.
+After obtaining the grasp groups, we perform triplet combinations of the groups to identify three groups that satisfy the approximately orthogonal constraint.
 To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triplet combination.
@@
 By performing successive regrasp using the prioritized triplets, object uncertainties can be effectively minimized.
-Although we consistently refer to “orthogonal” grasps or groups, it is important to note that strict 90-degree orthogonality is not required to determine the object’s pose.
+It is important to note that strict 90-degree orthogonality is not required to determine the object’s pose.
 The example shown in Fig.
@@
 We will discuss this aspect in more detail in the admittance control section.
-After identifying orthogonal groups, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups.
+After identifying the approximately orthogonal groups, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups.
 We use an incremental method to explore combinations of grasp triplets from the orthogonal groups and output a successful motion once found.
@@
 In this way, we incrementally explore the combinations of the grasp triplets and output successful motion once found.
-The detailed algorithmic flow is illustrated in Fig.
+The detailed algorithmic flow is shown in Fig.
 3.
@@
 If a large amount of data can be accumulated using the proposed method in the future, we are interested in exploring probabilistic modeling methods to learn grasp triplets and thus improve planning efficiency.
-However, this is beyond the scope of the present paper and is therefore not elaborated here.
+However, this is beyond the current scope and will not be elaborated.
 IV Conforming to Object Pose Uncertainty
@@
 Figure 8: (a) L-shape object and its random initial placements. (b) TL-shape object and its random initial placements.
-Table I summarizes the results.
+Table I shows the results.
 The data section has two main column groups reporting the positional differences and rotational differences.
 Within each group, there are three subcolumns.
-The $g_{2}$ columns show the positional and rotational differences of the $g_{2}$ grasp before and after admittance control.
-The $g_{3}$ columns present the corresponding differences for the $g_{3}$ grasp.
+The “second grasp” columns show the positional and rotational differences of the $g_{2}$ grasp before and after admittance control.
+The “third grasp” columns present the corresponding differences for the $g_{3}$ grasp.
 The “object” columns report the differences between the object pose estimated by the proposed method and the pose measured by the optical tracking system.
-The results for $g_{1}$ are not included, as it involves no admittance control.
+The “first grasp” uses position control and is thus excluded.
 TABLE I: Experimental Results with Statistics
-Task Position Differences ($\Delta\mathbf{p}$, mm) Rotation Vector Differences ($\Delta\mathbf{w}$, degrees)
-$g_{2}$ $g_{3}$ object $g_{2}$ $g_{3}$ object
+Position Differences ($\Delta\mathbf{p}$, mm) Rotation Vector Differences ($\Delta\mathbf{w}$, degrees)
+second grasp third grasp object second grasp third grasp object
 L1 -0.40 -1.60 5.20 -8.90 -2.00 -0.80 -33.70 -61.80 32.20 -0.55 -0.91 -0.02 -0.34 -0.75 -0.91 -0.01 2.82 -1.39
@@
 These results suggest that while the overall estimation errors appear large, their internal consistency is high, indicating that the system’s inaccuracies stem largely from systematic biases.
-Likely sources include errors from the optical tracking system itself (due to the 3D-printed marker mounts for calibration and robot frame location), installation errors in the dual-arm robot, and absolute control errors of the robot.
+Likely sources include calibration errors in the optical tracking system, installation errors in the two robot arms and the mounting aluminum frame, and absolute control errors of the robots.
 In addition, when comparing the deviations between the L-shaped and TL-shaped objects, we found that the inference results for the L-shaped object were better than those for the TL-shaped object.
 This outcome differs from our theoretical expectations.
-Theoretically, even if the three grasps are not perfectly orthogonal, the final inference accuracy should not degrade, and the L- and TL-shaped objects should exhibit similar deviations.
-We believe that the observed difference arises because the TL-shaped object involves rotations around axes unaligned with body frames, which place higher demands on the assembly precision of the dual-arm robot.
-The larger deviations reveal that the robot setup has greater misalignment in rotational components.
+We can conclude from the equations in Section V that even if the three grasps are not perfectly orthogonal, the final inference accuracy should not degrade, and the L- and TL-shaped objects should exhibit similar deviations.
+We believe that the observed difference arises because the TL-shaped object involves rotations around axes unaligned with body frames.
+The large difference indicates that the robot setup has greater rotational misalignment.
+This raises the question of compensating for system bias to improve absolute accuracy.
+While theoretically feasible, such correction would require complex estimation of both positional and rotational errors, which we leave for future work.
 In conclusion, although the results exhibit large absolute errors, the deviation levels are on the same order of magnitude as the optical tracking system.
@@
 The results demonstrate that the proposed approach can effectively reduce uncertainties and achieves high repeatability accuracy.
-The current method assumes planar contacts between flat fingertips and flat object surfaces.
-When the object surface is curved, the number of degrees of freedom constrained by a grasp will vary, and more grasps may be required to constrain uncertainties.
-Investigating the minimal and sufficient set of grasps under such conditions is an interesting future direction.
-Additionally, exploring learning-based methods to predict grasps and motion will be helpful to improve efficiency.
+In our experiments, we used a dual-arm setup built from two single-arm robots mounted on an aluminum frame, which inevitably introduced systematic errors and led to relatively large absolute errors.
+As a next step, we plan to validate the method in a dynamic simulation environment, where no absolute error is present, and use a dual-arm robot with commercially-calibrated body integration for practical deployment.